OpenStudy (anonymous):

If a factory that manufactures light bulbs, 0.04% of the bulbs manufactured are defective. It is expected that there will be one defective light bulb in what number of bulbs that are manufactured? a) 2500 b) 1250 c) 1000 d) 500 e) 250

OpenStudy (anonymous):

\[ 0.04\%\, n = 1 \]

OpenStudy (anonymous):

250 I think would be the answer. @wio

OpenStudy (e.mccormick):

Not 250. Percent means per centum, or per hundred. So one percent means one in a hundred. Four percent means four in a hundred... but you don't have four percent, you have point zero four percent... so four in how many? Then simplify that.

OpenStudy (anonymous):

My apologies but I don't understand quite well how to set this up. 4 in I'm not sure how many.

OpenStudy (e.mccormick):

OK. So how would you get rid of the decimal? Remember that per also means it is a fraction.

OpenStudy (anonymous):

wait, would it be 2,500?

OpenStudy (e.mccormick):

Yes. \(\dfrac{.04}{100}\cdot \dfrac{100}{100}=\dfrac{4}{10000}\) so 4 in 10,000 or one in 2,500.

OpenStudy (anonymous):

great! Thank you very much for your help! Would you mind helping me on another problem please? If not, it's fine. :)

OpenStudy (e.mccormick):

What is it?

OpenStudy (anonymous):

Its another word problem like this one.

OpenStudy (e.mccormick):

Not type... the actual problem. =)

OpenStudy (anonymous):

Three students use a computer for a total of 3 hours. If the first student uses the computer 28% of the total time, and the second student uses the computer 52% of the total time, how many minutes does the third student use the computer? a) 24 b) 30 c) 36 d) 42 e) 50

OpenStudy (e.mccormick):

OK. First, pull out the numbers and try to see how they relate to the answer you need. So, what important numbers do you have?

OpenStudy (anonymous):

3 hours, 28% and 52%

OpenStudy (e.mccormick):

Now, you have 3 hours... but what is the answer asking for?

OpenStudy (anonymous):

hold for one second please. :)

OpenStudy (anonymous):

ok, I am back. Thank you for waiting.

OpenStudy (anonymous):

It is asking how much time the third student used the computer for.

OpenStudy (e.mccormick):

But time in what units?

OpenStudy (anonymous):

minutes.

OpenStudy (e.mccormick):

So, you have hours but need minutes. Might as well fix that right off.

OpenStudy (anonymous):

ok. I'm just a tid bit confused how to set it up.

OpenStudy (e.mccormick):

Well, the total percentage of all three is 100%, right? So, you are given the persentage for two of them. How will you go about finding the percentage of the third?

OpenStudy (anonymous):

100-80= 20?

OpenStudy (anonymous):

as in 20%...

OpenStudy (e.mccormick):

OK. So you need 20% of 3 hours in minutes.

OpenStudy (anonymous):

3 hours in minutes is 180, but I'm not sure we need to convert that. would 20% turn into .20 ?

OpenStudy (e.mccormick):

You do need to conver the hours, and yes, .2

OpenStudy (anonymous):

i keep getting 15 but that would be wrong since it is not on the answer choices..

OpenStudy (e.mccormick):

Seems you are divinding 3 by .2 But a percentage of in this format is multiplied... and remember, you need minutes!

OpenStudy (anonymous):

...I honestly don't know. :/ my apologies.

OpenStudy (e.mccormick):

Well, look at this another way. A dollar! 25% of one dolar is .25 * one dollar... or .25 dollars. But we have another word for that. Cents. So 25 cents. Can you do the same vasic thing with your problem?

OpenStudy (anonymous):

24 minutes?

OpenStudy (e.mccormick):

Hmm... what did you multiply to get that?

OpenStudy (anonymous):

actally, im completly wrong. I dont know what I did. I need to go do something'very important quickly. I dont know if you will still be on...but I will keep working on this problem.

OpenStudy (e.mccormick):

Let me try a different example. Say I needed to measure 13% of two miles in feet. So, I need to set this up. Well, one mile is 5280 feet. 13% is .13. So I put the whole thing in one equation: \(2\times 5280\times .13\) The 2 is the 2 miles. The 5280 is to get the answer in feet. The .13 is the 13%. \(2\times 5280\times .13 = 1372.8\) feet. You are doing this same thing! Only the numbers you plug in are different. The basic equation is the same.